Consider the following Keynesian income model:

E = C + I + G + X-M

C = 300 + 0.85Yd

Yd = Y – T

T = 60 + 0.25Y; I = 400

G = 700

X = 400

M = 50 + 0.15Y

In equilibrium, Y = E:

a. calculate the equilibrium level of income.

b. calculate the amount of taxes collected when the economy is at equilibrium level of income and show whether the government budget is in surplus or deficit.

c. calculate the value of net exports when the economy is at equilibrium level of income.
Answer #1

At Equilibrium > y = c+I + G + X-M + y = 300 t 0.851d + 400+ Toof 400- (50 fo15Y] → Ya 1800 +0.8517- 60 - -25Y) - 50-3157 Y =- value of Net Exports be M=507.154 am=50 f 15/33 15.12) */M = 547-268] The value of Net Exhauts when the economy is at Equil